# Dynamic 3D Gaussian Fields for Urban Areas

### v1

@@ 1 Introduction
The problem of synthesizing novel views from a set of images has received widespread attention in recent years due to its importance for technologies like AR/VR and robotics. In particular, obtaining interactive, high-quality renderings of large-scale, dynamic urban areas under varying weather, lightning, and seasonal conditions is a key requirement for closed-loop robotic simulation and immersive VR experiences. To achieve this goal, sensor-equipped vehicles act as a frequent data source that is becoming widely available in city-scale mapping and autonomous driving, creating the possibility of building up-to-date digital twins of entire cities. However, modeling these scenarios is extremely challenging as heterogeneous data sources have to be processed and combined: different weather, lighting, seasons, and distinct dynamic and transient objects pose significant challenges to the reconstruction and rendering of dynamic urban areas.
@@ 2 Related Work
Reconstructing urban areas. Dynamic urban areas are particularly challenging to reconstruct due to the complexity of both the scenes and the capturing process. Hence, significant research efforts have focused on adapting view synthesis approaches from controlled, small-scale environments to larger, real-world scenes. In particular, researchers have investigated the use of depth priors from e.g. LiDAR, providing additional information such as camera exposure, jointly optimizing camera parameters, and developing specialized sky and light modeling approaches [8, 9, 10, 11, 12, 13, 14]. However, since scene dynamics are challenging to approach, many works simply remove dynamic areas, providing only a partial reconstruction. Fewer works explicitly model scene dynamics [16, 15, 49, 45] and consequently, these works have clear limitations: they do not scale beyond single, short video clip [15, 49, 45], struggle to accurately represent dynamic objects [16], or are costly to render [17, 76]. Instead, we present a fast, scalable method that faithfully represents scene dynamics.
@@ 3.2 Representation
Figure 2: Overview. We use sets of 3D Gaussians $G$ as geometry scaffold, neural fields $\phi$ and $\psi$ to represent sequence- and object-specific appearance and geometry variations, and a scene graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ to express the scene configuration at each sequence-time pair $(s,t)$. We condition the neural fields with latent codes $\omega$ of the nodes in $\mathcal{V}$. To render a view at $(s,t)$, we compose the sets of 3D Gaussians using the coordinate system transformations $[\mathbf{R}|\mathbf{t}]$ along the edges $\mathcal{E}$.
@@
Differently from [18], our 3D Gaussians do not hold any appearance information, reducing the memory footprint of the representation by more than 80%. Instead, we leverage neural networks to regress a color information $\mathbf{c}_{k}^{s,t}$ and an updated opacity $\alpha_{k}^{s,t}$ for each sequence $s\in S$ and time $t\in T_{s}$. For 3D Gaussians in $G_{r}$ modeling the scene scaffolding, we predict an opacity attenuation term $\nu_{k}^{s,t}$ that is used to model transient geometry by downscaling $\alpha_{k}$ Instead, for 3D Gaussians in $G_{o}$ modeling objects the base opacity is left invariant. Hence
@@
Pruning decisions in ADC are obtained by thresholding the base opacity $\alpha_{k}$, which is directly accessible without computational overhead.
Finally, in the presence of non-rigid objects $o$, we predict deformation terms $\delta\boldsymbol{\mu}_{k}^{t}$ to the position of 3D primitives in $G_{o}$ via a neural network, for each time $t\in T_{o}$. In this case, the final position of the primitive is given by
$\boldsymbol{\mu}_{k}^{t}\coloneqq\boldsymbol{\mu}_{k}+\delta\boldsymbol{\mu}_{k}^{t}\,.$ (5)
Appearance and transient geometry. Given the scene graph $\mathcal{G}$ and the 3D Gaussians $G$, we use two efficient neural fields [53] to decode the appearance parameters of each primitive. For 3D Gaussians in $G_{r}$ modeling the static scene, the neural field is denoted by $\phi$ and regresses the opacity attenuation term $\nu_{k}^{s,t}$ and a color $\mathbf{c}_{k}$, given the 3D Gaussian primitive’s position $\boldsymbol{\mu}_{k}$, a viewing direction $\mathbf{d}$, the base opacity $\alpha_{k}$ and the latent code of the node $\omega_{s}^{t}$, i.e.
@@
where $s\in S$ and $t\in T_{s}$. The opacity attenuation contributes to modeling transient geometry, for it potentially enables removing parts of the scene encoded in the original set of Gaussians. Moreover, it does not depend on the viewing direction $\mathbf{d}$.
For 3D Gaussians in $G_{o}$ modeling dynamic objects, the neural field is denoted by $\psi$ and regresses a color $\mathbf{c}^{t}_{k}$. Besides the primitive’s position and viewing direction, we condition $\psi$ on latent vectors $\omega_{s}^{t}$ and $\omega_{o}^{t}$ to model both local object texture and global sequence appearance such as illumination. Here, the sequence $s$ is the one where $o$ belongs to, i.e. satisfying $o\in O_{s}$, and $t\in T_{o}$. Accordingly, the color $\mathbf{c}^{t}_{k}$ for a 3D Gaussian in $G_{o}$ is given by
@@
Both $\boldsymbol{\mu}_{k}$ and $\mathbf{d}$ are expressed in the canonical, object-centric space of object $o$.
Using neural fields has three key advantages for our purpose. First, by sharing the parameters of $\phi$ and $\psi$ across all 3D Gaussians $G$, we achieve a significantly more compact representation than in [18] when scaling to large-scale urban areas. Second, it allows us to model sequence-dependent appearance and transient geometry which is fundamental to learning a scene representation from heterogeneous data. Third, information sharing between nodes enables an interaction of sequence and object appearance.
Non-rigid objects. Street scenes are occupied not only by rigidly moving vehicles but also by, e.g., pedestrians and cyclists that move in a non-rigid manner. These pose a significant challenge due to their unconstrained motion under limited visual coverage. Therefore, we take a decomposed approach to modeling non-rigid objects that uses the scene graph $\mathcal{G}$ to model the global, rigid object motion while using a deformation head $\chi$ to model the local, articulated motion. The deformation head predicts a local position offset $\delta\boldsymbol{\mu}_{k}^{t}$ via
$\delta\boldsymbol{\mu}_{k}^{t}\coloneqq\chi(\mathbf{f}_{\psi},\gamma(t))$ (8)
given an intermediate feature representation $\mathbf{f}_{\psi}$ of $\psi$ conditioned on $\boldsymbol{\mu}_{k}$. This way, we can deform the position of $\boldsymbol{\mu}_{k}$ over time in canonical space as per Equation 5.
@@ 3.3 Composition and Rendering
Scene composition. To render our representation from the perspective of camera $c$ at time $t$ in sequence $s$, we traverse the graph $\mathcal{G}$ to obtain the latent vector $\omega_{s}^{t}$ and the latent vector $\omega_{o}^{t}$ of each visible object $o\in O_{s}$, i.e. such that $t\in T_{o}$. Moreover, for each 3D Gaussian primitive $\mathfrak{g}_{k}$ in $G$, we use the collected camera parameters, object scale, and pose information to determine the transformation $\Pi^{c}_{k}$ mapping points from the primitive’s reference frame (e.g. world for $G_{r}$, object-space for $G_{o}$) to the image space of camera $c$. Opacities $\alpha^{s,t}_{k}$ are computed as per Equation 4, while colors $\mathbf{c}^{s,t}_{k}$ are computed for primitives in $G_{r}$ and in $G_{o}$ via Equations 6 and 7, respectively.
@@ 3.4 Optimization
To optimize parameters $\theta$ of $f_{\theta}$, i.e. 3D Gaussian parameters $\mu_{k}$, $\alpha_{k}$, $q_{k}$ and $a_{k}$, sequence latent vectors $\omega_{s}^{t}$ and implicit neural fields $\psi$ and $\phi$, we use an end-to-end differentiable rendering pipeline. We render both an RGB color image $\hat{\mathtt{I}}$ and a depth image $\hat{\mathcal{D}}$ and apply the following loss function:
@@
Adaptive density control. To facilitate the growth and pruning of 3D Gaussian primitives, the optimization of the parameters $\theta$ is interleaved by an ADC mechanism [18]. This mechanism is essential to achieve photo-realistic rendering. However, it was not designed for training on tens of thousands of images, and thus we develop a streamlined multi-GPU variant. First, accumulating statistics across processes is essential. Then, instead of running ADC on GPU 0 and synchronizing the results, we synchronize only non-deterministic parts of ADC, i.e. the random samples drawn from the 3D Gaussians that are being split. These are usually much fewer than the total number of 3D Gaussians and thus avoids communication overhead. Next, the 3D Gaussian parameters are replaced by their updated replicas. However, this will impair the synchronization of the gradients because, in PyTorch DDP [79], parameters are only registered once at model initialization. Therefore, we re-initialize the Reducer upon finishing the ADC mechanism in the low-level API provided in [79].
@@ 4 Experiments
Datasets and Metrics. We evaluate our approach across three dynamic outdoor benchmarks. First, we utilize the recently proposed NVS benchmark [17] on Argoverse 2 [80] to compare against the state-of-the-art in the multi-sequence scenario and to showcase the scalability of our method. Second, we use the established Waymo Open [23], KITTI [21] and VKITTI2 [22] benchmarks to compare to existing approaches in single-sequence scenarios. For Waymo, we use the dynamic-32 split of [76], while for KITTI and VKITTI2 we follow [16, 17]. We apply commonly used metrics to measure view synthesis quality: PSNR, SSIM [78], and LPIPS (AlexNet) [81].
@@ 4.1 Comparison to State-of-the-Art
Multi-sequence setting. In Table 1, we show results on the Argoverse 2 NVS benchmark proposed in [17]. We compare to state-of-the-art approaches [16, 17] and the baselines introduced in [17]. The results highlight that our approach scales well to large-scale dynamic urban scenes, outperforming previous work in performance and rendering speed by a significant margin. Specifically, we outperform [17] by more than 3 points in PSNR while rendering more than 200×200$\times$200 × faster. To examine these results more closely, we show a qualitative comparison in Figure 3. We see that while SUDS [16] struggles with dynamic objects and ML-NSG [17] produces blurry renderings, our work provides sharp renderings and accurately represented dynamic objects, in both RGB color and depth images. Overall, the results highlight that our model can faithfully represent heterogeneous data at high visual quality in a single 3D representation while being much faster to render than previous work.
Single-sequence setting. In Table 2, we show results on the KITTI [21] and VKITTI [22] benchmarks at varying training split fractions. Our approach outperforms previous work as well as concurrent 3D Gaussian-based approaches [73]. In Table 3, we show results on Waymo Open [23], specifically on the Dynamic-32 split proposed in [76]. We outperform previous work by a large margin while our rendering speed is 700×700$\times$700 × faster than the best alternative. Note that our rendering speed increases for smaller-scale scenes. Furthermore, we show that, contrary to previous approaches, our method does not suffer from lower view quality in dynamic areas. This corroborates the strength of our contributions, showing that our method is not only scalable to large-scale, heterogeneous street data but also demonstrates superior performance in smaller-scale, homogeneous street data.
Methods Full Image Dynamic-Only Render (s)

### v2

@@ 1 Introduction
The problem of synthesizing novel views from a set of images has received widespread attention in recent years due to its importance for technologies like AR/VR and robotics. In particular, obtaining interactive, high-quality renderings of large-scale, dynamic urban areas under varying weather, lighting, and seasonal conditions is a key requirement for closed-loop robotic simulation and immersive VR experiences. To achieve this goal, sensor-equipped vehicles act as a frequent data source that is becoming widely available in city-scale mapping and autonomous driving, creating the possibility of building up-to-date digital twins of entire cities. However, modeling these scenarios is extremely challenging as heterogeneous data sources have to be processed and combined: different weather, lighting, seasons, and distinct dynamic and transient objects pose significant challenges to the reconstruction and rendering of dynamic urban areas.
@@ 2 Related Work
Reconstructing urban areas. Dynamic urban areas are particularly challenging to reconstruct due to the complexity of both the scenes and the capturing process. Hence, significant research efforts have focused on adapting view synthesis approaches from controlled, small-scale environments to larger, real-world scenes. In particular, researchers have investigated the use of depth priors from e.g. LiDAR, providing additional information such as camera exposure, jointly optimizing camera parameters, and developing specialized sky and light modeling approaches [8, 9, 10, 11, 12, 13, 14]. However, since scene dynamics are challenging to approach, many works simply remove dynamic areas, providing only a partial reconstruction. A few works explicitly model scene dynamics, but suffer from limitations in terms of scalability [15, 49, 45], accuracy [16], rendering speed [17, 76, 77], or modeling of non-rigid and uncommon objects [15, 49, 45, 77, 17]. We introduce a mechanism to handle transient geometry and varying appearance, improve rendering efficiency, and, inspired by how global rigid object motion is handled in [17, 15, 49, 45, 77], propose an approach to model the local articulated motion of non-rigid dynamic objects without using semantic priors. Consequently, our work enables the reconstruction of much larger urban areas with a significantly higher number and diversity of dynamic objects across multiple in-the-wild captures.
@@ 3.2 Representation
Figure 2: Overview. To render an image of sequence $s$ at time $t$, we first evaluate the scene graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ which stores latent codes $\omega$ at its nodes $\mathcal{V}$ and coordinate transformations $[\mathbf{R}|\mathbf{t}]$ at its edges $\mathcal{E}$, i.e. the configuration of the dynamic objects and the overall scene. We then use the scene configuration to determine the active sets of 3D Gaussians $G$. The 3D Gaussians $G$ and the latent codes $\omega$ serve as conditioning signals to the neural fields $\phi$ and $\psi$, which output, for each 3D Gaussian $\mathfrak{g}_{k}\in G$, an appearance conditioned color $\mathbf{c}^{s,t}_{k}$, an opacity correction term $\nu^{s,t}_{k}$ for static Gaussians modeling transient geometry, and a dynamic deformation $\delta^{t}_{k}$ for non-rigid dynamic 3D Gaussians modeling e.g. pedestrians. Finally, the retrieved information is used to compose a set of 3D Gaussians that represent the dynamic scene at $(s,t)$ from which we render the image.
@@
Differently from [18], our 3D Gaussians do not hold any appearance information, reducing the memory footprint of the representation by more than 80%. Instead, we leverage neural fields to regress a color information $\mathbf{c}_{k}^{s,t}$ and an updated opacity $\alpha_{k}^{s,t}$ for each sequence $s\in S$ and time $t\in T_{s}$. For 3D Gaussians in $G_{r}$ modeling the scene scaffolding, we predict an opacity attenuation term $\nu_{k}^{s,t}$ that is used to model transient geometry by downscaling $\alpha_{k}$. Instead, for 3D Gaussians in $G_{o}$ modeling objects the base opacity is left invariant. Hence
@@
The attenuation term enforces a high base opacity for every 3D Gaussian visible in at least one sequence. Therefore, we can obtain pruning decisions in ADC by thresholding the base opacity $\alpha_{k}$, which is directly accessible without computational overhead, without risking the removal of transient geometry.
Furthermore, in the presence of non-rigid objects $o$, we predict deformation terms $\delta_{k}^{t}\in^{3}$ to the position of 3D primitives in $G_{o}$ via the neural fields, for each time $t\in T_{o}$. In this case, the position of the primitive in object-centric space at time $t$ is given by
$\boldsymbol{\mu}_{k}^{t}\coloneqq\boldsymbol{\mu}_{k}+\delta_{k}^{t}\,.$ (5)
Appearance and transient geometry. Given the scene graph $\mathcal{G}$ and the 3D Gaussians $G$, we use two neural fields to decode the aforementioned parameters for each primitive. In particular, for 3D Gaussians in $G_{r}$ modeling the static scene, the neural field is denoted by $\phi$ and regresses the opacity attenuation term $\nu_{k}^{s,t}$ and a color $\mathbf{c}^{s,t}_{k}$, given the 3D Gaussian primitive’s position $\boldsymbol{\mu}_{k}$, a viewing direction $\mathbf{d}$, the base opacity $\alpha_{k}$ and the latent code of the node $\omega_{s}^{t}$, i.e.
@@
where $s\in S$ and $t\in T_{s}$. Note that since the opacity attenuation $\nu_{k}^{s,t}$ contributes to modeling transient geometry by removing parts of the scene encoded in the original set of Gaussians, it does not depend on the viewing direction $\mathbf{d}$.
For 3D Gaussians in $G_{o}$ modeling dynamic objects, the neural field is denoted by $\psi$ and regresses a color $\mathbf{c}^{s,t}_{k}$. Besides the primitive’s position and viewing direction, we condition $\psi$ on latent vectors $\omega_{s}^{t}$ and $\omega_{o}^{t}$ to model both local object texture and global sequence appearance such as illumination. Here, the sequence $s$ is the one where $o$ belongs to, i.e. satisfying $o\in O_{s}$, and $t\in T_{o}$. Accordingly, the color $\mathbf{c}^{s,t}_{k}$ for a 3D Gaussian in $G_{o}$ is given by
@@
Both $\boldsymbol{\mu}_{k}$ and $\mathbf{d}$ are expressed in the canonical, object-centric space of object $o$. Using neural fields has three key advantages for our purpose. First, by sharing the parameters of $\phi$ and $\psi$ across all 3D Gaussians $G$, we achieve a significantly more compact representation than in [18] when scaling to large-scale urban areas. Second, it allows us to model sequence-dependent appearance and transient geometry which is fundamental to learning a scene representation from heterogeneous data. Third, information sharing between nodes enables an interaction of sequence and object appearance.
However, querying a neural field is more complex than a spherical harmonics function as in [18]. Therefore, we i) use efficient hash-grid representations [53] to minimize query complexity and, ii) carefully optimize the rendering workflow to minimize the amount of queries. In particular, we skip out-of-view 3D Gaussians and implement a vectorized query function to $\psi$ that retrieves the parameters of all relevant dynamic objects in parallel. We refer to Section 4.3 for a runtime analysis.
Non-rigid objects. Street scenes are occupied not only by rigidly moving vehicles but also by, e.g., pedestrians and cyclists that move in a non-rigid manner. These pose a significant challenge due to their unconstrained motion under limited visual coverage. Therefore, we take a decomposed approach to modeling non-rigid objects. First, we represent the local articulated motion of non-rigid objects like pedestrians as deformation in canonical space. We use deformation head $\chi$ that predicts a local position offset $\delta_{k}^{t}$ via
$\delta_{k}^{t}\coloneqq\chi(\mathbf{f}_{\psi},\gamma(t))$ (8)
given an intermediate feature representation $\mathbf{f}_{\psi}$ of $\psi$ conditioned on $\boldsymbol{\mu}_{k}$ and time $t$. We deform the position of $\boldsymbol{\mu}_{k}$ over time in canonical space as per Equation 5. Second, we use the scene graph $\mathcal{G}$ to model the global rigid object motion, transforming the objects from object-centric to world space with a rigid body transformation. We use a general design to cover a wide range of scenarios, such as pedestrians holding a stroller or shopping bags, cyclists, and animals. See Figure 10 in our supp. mat.
@@ 3.3 Composition and Rendering
Scene composition. To render our representation from the perspective of camera $c$ at time $t$ in sequence $s$, we traverse the graph $\mathcal{G}$ to obtain the latent vector $\omega_{s}^{t}$ and the latent vector $\omega_{o}^{t}$ of each visible object $o\in O_{s}$, i.e. such that $t\in T_{o}$. Moreover, for each 3D Gaussian primitive $\mathfrak{g}_{k}$ in $G$, we use the collected camera parameters, object scale, and pose information to determine the transformation $\Pi^{c}_{k}$ mapping points from the primitive’s reference frame (e.g. world for $G_{r}$, object-space for $G_{o}$) to the image space of camera $c$. Opacities $\alpha^{s,t}_{k}$ are computed as per Equation 4, while colors $\mathbf{c}^{s,t}_{k}$ are computed for primitives in $G_{r}$ and in $G_{o}$ via Equations 6 and 7, respectively. For non-rigid objects in $G_{o}$, we compute the primitive positions $\boldsymbol{\mu}^{t}_{k}$ via Equation 5.
@@ 3.4 Optimization
To optimize parameters $\theta$ of $f_{\theta}$, i.e. 3D Gaussian parameters $\boldsymbol{\mu}_{k}$, $\alpha_{k}$, $q_{k}$ and $a_{k}$, sequence latent vectors $\omega_{s}^{t}$ and implicit neural fields $\psi$ and $\phi$, we use an end-to-end differentiable rendering pipeline. We render both an RGB color image $\hat{\mathtt{I}}$ and a depth image $\hat{\mathcal{D}}$ and apply the following loss function:
@@
Adaptive density control. To facilitate the growth and pruning of 3D Gaussian primitives, the optimization of the parameters $\theta$ is interleaved by an ADC mechanism [18]. This mechanism is essential to achieve photo-realistic rendering. However, it was not designed for training on tens of thousands of images, and thus we develop a streamlined multi-GPU version of it. We accumulate statistics across processes and, instead of running ADC on GPU 0 and synchronizing the results, we synchronize only non-deterministic parts of ADC, i.e. the random samples drawn from the 3D Gaussians that are being split. These are usually much fewer than the total number of 3D Gaussians and thus avoids communication overhead. Next, the 3D Gaussian parameters are replaced by their updated replicas. However, this will impair the synchronization of the gradients because, in PyTorch DDP [80], parameters are only registered once at model initialization. Therefore, we re-initialize the Reducer upon finishing the ADC mechanism in the low-level API provided in [80].
@@ 4 Experiments
Datasets and metrics. We evaluate our approach across various dynamic outdoor benchmarks. First, we utilize the recently proposed NVS benchmark [17] of Argoverse 2 [81] to compare against the state-of-the-art in the multi-sequence scenario and to showcase the scalability of our method. Second, we use the established Waymo Open [23], KITTI [21] and VKITTI2 [22] benchmarks to compare to existing approaches in single-sequence scenarios. For Waymo, we use the dynamic-32 split of [76], while for KITTI and VKITTI2 we follow [16, 17]. We apply commonly used metrics to measure view synthesis quality: PSNR, SSIM [79], and LPIPS (AlexNet) [82].
@@ 4.1 Comparison to state-of-the-art
Multi-sequence setting. In Table 1, we show results on the Argoverse 2 NVS benchmark proposed in [17]. We compare to state-of-the-art approaches [16, 17] and the baselines introduced in [17]. The results highlight that our approach scales well to large-scale dynamic urban scenes, outperforming previous work in performance and rendering speed by a significant margin. Specifically, we outperform [17] by more than 3 dB in PSNR while rendering more than 200×200$\times$200 × faster. To examine these results more closely, we show a qualitative comparison in Figure 3. We see that while SUDS [16] struggles with dynamic objects and ML-NSG [17] produces blurry renderings, our work provides sharp renderings and accurately represented dynamic objects, in both RGB color and depth images. Overall, the results highlight that our model can faithfully represent heterogeneous data at high visual quality in a single 3D representation while being much faster to render than previous work.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
SUDS [16, 83] 23.12 0.821 0.135
MARS [83] 24.00 0.801 0.164
NeuRAD [77] 27.00 0.795 0.082
NeuRAD-2x [77] 27.91 0.822 0.066
4DGF (Ours) 30.01 0.913 0.052
Table 3: Novel view synthesis on KITTI [21] on split from [77]. We follow the experimental protocol in [77] to compare to additional works. baseline from [83].
Single-sequence setting. In Table 2, we show results on the KITTI [21] and VKITTI [22] benchmarks at varying training view fractions, i.e. from dense towards sparse view setting. Furthermore, we follow the experimental protocol in [77] and show an additional comparison on the KITTI dataset with a different data split that uses also approx. 50% of views for training in Table 3. Our approach consistently outperforms previous work as well as concurrent 3D Gaussian-based approaches [73]. In Table 4, we show results on Waymo Open [23], specifically on the Dynamic-32 split proposed in [76]. We outperform previous work by a large margin while our rendering speed is 700×700$\times$700 × faster than the best alternative. Note that our rendering speed increases for smaller-scale scenes. Furthermore, we show that, contrary to previous approaches, our method does not suffer from lower view quality in dynamic areas. This corroborates the strength of our contributions, showing that our method is not only scalable to large-scale, heterogeneous street data but also demonstrates superior performance in smaller-scale, homogeneous street data.
Method Full Image Dynamic-Only Render (s)
@@
4.2 Ablation studies
We verify our design choices in both multi- and single-sequence settings. For a fair comparison, we set the global maximum of 3D Gaussians to 8 and 4.1 million, respectively. We perform these ablation studies on the residential split of [17]. We use the full overlap in the multi-sequence setting, while using a single sequence of this split for the single-sequence setting. In Table 5a, we verify the components that are not specific to the multi-sequence setting. In particular, we show that our approach to modeling scene dynamics is highly effective, evident from the large disparity in performance between the static and the dynamic variants. Next, we show that modeling appearance with a neural field is on par with the proposed solution in [18], while being more memory efficient. In particular, when modeling view-dependent color as a per-Gaussian attribute as in [18] the model uses 8.6 GB of peak GPU memory during training, while it uses only 4.5 GB with fixed-size neural fields. Similarly, storing the parameters of the former takes 922 MB, while the latter takes only 203 MB. Note that this disparity increases with the number of 3D Gaussians per scene. Finally, we achieve the best performance when adding the generated 3D Gaussian background.
4.3 Runtime analysis
We divide our algorithm into its main components and report the individual inference runtimes across two datasets in Table 6. While the runtime clearly correlates with scene complexity and image resolution, we observe that, on average, the runtime is dominated by scene graph evaluation and rasterization, accounting for more than 75% of the total runtime. This owes to the complexity of rasterizing millions of primitives across a high-resolution image which is computationally demanding even for efficient rasterization algorithms [18], and handling hundreds to thousands of dynamic objects across one or multiple dynamic captures, making the retrieval of the 3D Gaussians and latent codes costly. In contrast, the queries to the neural fields account for only 13.3% of the average total runtime, making it a viable alternative to the spherical harmonics function in [18]. Overall, our method achieves interactive rendering speeds on both datasets and 20.4 FPS on average.